Tom and Tim ages add up to 21 years. In 3 years Tom will be twice tim's age. How old are they now?

Answer by ewatrrr(24785)

 
Hi
Let x represent tom's age and y Tim's age
x + y = 21 0r y = 21-x
In 3 years...
x + 3 = 2[(21-x)+3]
x + 3 = 2[24-x]
x + 3 = 48 - 2x
3x = 45
x = 15, Tom's age. Tim is 6 21-15
CHECKING our answer***In 3 years
18 = 2•9
Wish You the Best in your Studies.